Find the two roots of (x-3)^2 = 25 without changing (x-3)^2 to a trinomial.
Answer by stanbon(75887)

Find the two roots of (x-3)^2 = 25
(x-3)^2 - 25 = 0
Factor to get:
(x-3 - 5)(x-3 + 5) = 0
(x-8)(x+2) = 0
x = 8 or x = -2
==================
Cheers,
Stan H.